# A characterization of the local structure of two-dimensional sets with positive reach

### v1

@@ 4. Local structure of $T_{k}(A)$
where (13) and the $L$-Lipschitzness of $\psi_{k}$ on $M$ were used, so (2) and (3) follow with $c:=\max\{L,\frac{1}{2r}\}$. Now, using Fact 2.3, there exists a $C^{1,1}$ mapping $F:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $F(x)=f(x)$ and $DF(x)=\varphi(x)$ whenever $x\in M$. Note that $F(0)=f(0)=0$. Since $DF(0)=\varphi(0)$ is the identity map, we can apply Fact 2.2 to $F$ and find $\omega>\varepsilon>0$ such that the restriction $\Phi|_{B(0,\varepsilon)}$ is a $C^{1,1}$-diffeomorphism. Choose $\varepsilon>\delta>0$ such that $B(0,\delta)\subset\Phi(B(0,\varepsilon))$ and set $G(x):=\pi_{T^{\perp}}(\Phi^{-1}(x))$ for $x\in T\cap B(0,\delta)$. Then $G:T\cap B(0,\delta)\to T^{\perp}$ is a $C^{1,1}$ mapping which has a $C^{1,1}$ extension $\tilde{G}:T\to T^{\perp}$ by Fact 2.4. Set $\Gamma:=\{x+\tilde{G}(x):\ x\in T\}$. Then $\Gamma$ is a $k$-dimensional $C^{1,1}$ surface and $T_{k}(A)\cap B(a,\delta)\subset\Gamma$. Indeed, consider an arbitrary $z\in T_{k}(A)\cap B(a,\delta)\subset M$ and denote $x:=\Phi(z)=f(z)=\pi_{T}(z)$. Then $G(x)=\pi_{T^{\perp}}(\Phi^{-1}(x))=\pi_{T^{\perp}}(z)$ and so $z=\pi_{T}(z)+\pi_{T^{\perp}}(z)=x+G(x)\in\Gamma$. ∎
@@
By Theorem 4.1 we can assign to each $a\in T_{k}(A)$ a $\delta_{a}>0$ and a $k$-dimensional $C^{1,1}$ surface $\Gamma_{a}$ such that $T_{k}(A)\cap B(a,\delta_{a})\subset\Gamma_{a}$. Since $\mathbb{R}^{d}$ is a separable metric space, there exists a countable set $S\subset T_{k}(A)$ such that $\bigcup_{a\in T_{k}(A)}B(a,\delta_{a})=\bigcup_{a\in S}B(a,\delta_{a})$. Then
@@
This proves (i). Further, we obtain (ii) from (i) using (14) since each $i$-dimensional $C^{1,1}$ surface with $1\leq i\leq k$ is clearly contained in a $k$-dimensional $C^{1,1}$ surface, and $T_{0}(A)$ is countable. ∎
@@
By [23, Theorem 7.5], AA can be locally covered by finitely many DC surfaces of dimension $k$ (which are even semiconcave if k=d1k=d-1, see [23, Theorem 5.9]). These results, however, do not imply Corollary 4.2 (ii) where $C^{1,1}$ surfaces are used.
@@
Without any loss of generality we will suppose that $a=0$ and $e_{1}\in\operatorname{Tan}(A,0)$. Denote $W:=\operatorname{span}{e_{1}}$, $\pi:=\pi_{W}$ and $V:=W^{\perp}$.
@@
$\varphi(w):=\gamma\left(\left(G|_{U}\right)^{-1}(w)\right)-w\in V,\quad w\in G(U)\subset W,$
$\varphi:G(U)\subset W\to V$ is $C^{1,1}$ smooth and
@@
Choose $0<\delta<\omega$ such that $(-\delta,\delta)\subset G(U)$. Then, by Fact 2.4, $\varphi|_{(-\delta,\delta)}:(-\delta,\delta)\to V$ has an $C^{1,1}$ extension $\tilde{\varphi}:W\to V$, $\Gamma:=\{w+\tilde{\varphi}(w):\,w\in W\}$ is a $1$-dimensional $C^{1,1}$ surface and we have
@@ 5.1. Description of the strategy and some auxiliary notions
The difficulty of Theorem 1.3 consists in the proof of its first part (the proof of the second - “moreover” - part is easy). It is advantageous to reformulate and prove the first part using an equivalence relation $\approx$ from the following definition.
@@
Let $A\subset\mathbb{R}^{d}$ ($d\geq 3$) have positive reach, $\dim A\leq 2$ and $a\in T_{1}(A)$. Then there exists a $\operatorname{B}$-set $B$ and a multirotation $\rho$ associated with $T_{1}(B)$ such that $(A,a)\approx(\rho(B),0)$.
@@
    There exists a $\operatorname{B}$-set $B_{2}$ and a multirotation $\rho_{2}$ associated with $T_{1}(B_{2})$ such that $(A,a)\approx(\rho_{2}(B_{2}),0)$.
@@
Let $C,C^{\prime}\subset W:=\operatorname{span}\{e_{1}\}\subset\mathbb{R}^{d}$ ($d\geq 3$) be nonempty compact sets such that $C\cap[p,q]=C^{\prime}\cap[p,q]$ for some $p<q$. Let $\rho$ be a multirotation associated with $C$. Then there exists a multirotation $\rho^{\prime}$ associated with $C^{\prime}$ such that $\rho|_{\pi^{-1}[p,q]}=\rho^{\prime}|_{\pi^{-1}[p,q]}$.
@@
Let $\rho$ be determined by $2$-rotations ($R_{\lambda}$), where $\lambda$ are components of $W\setminus C$. We define $\rho^{\prime}$ as multirotation determined by $2$-rotations $(R^{\prime}_{\lambda^{\prime}})$, where $\lambda^{\prime}$ are components of $W\setminus C^{\prime}$. We set $R^{\prime}_{\lambda^{\prime}}:={\rm id}$ if $\lambda^{\prime}\cap[p,q]=\emptyset$ and $R^{\prime}_{\lambda^{\prime}}:=R_{\lambda}$ if $\lambda^{\prime}\cap[p,q]\neq\emptyset$, where $\lambda$ is the (clearly unique) component of $W\setminus C$ such that $\lambda\cap[p,q]=\lambda^{\prime}\cap[p,q]$. It is easy to verify that $\rho^{\prime}$ has the desired properties. ∎
@@
Throughout the section we will use the usual identifications $\mathbb{R}^{d}=\mathbb{R}\times\mathbb{R}^{d-1}$ and $\mathbb{R}=\mathbb{R}\times\{0\}=:W\subset\mathbb{R}^{d}$. Further, we denote $\pi:=\pi_{W}$, $\pi:\mathbb{R}^{d}\to\mathbb{R}$.
@@
    Any $\operatorname{B}$-set $B\subset\mathbb{R}^{2}\subset\mathbb{R}^{d}$ is skewered (on $[0,r]$ or [r,r][-r,r] for some $r>0$), both in $\mathbb{R}^{2}$ and in $\mathbb{R}^{d}$. Indeed, $B$ has positive reach in $\mathbb{R}^{2}$ by Lemma 2.22, and hence also in $\mathbb{R}^{d}$ by Remark 2.14, and it is easy to see that $0\in T_{1}(B)$. Moreover, if $\varphi,\psi$ are the functions from Definition 1.1, we have
@@
The proof of Proposition 5.3 for the case $a\in T_{1}^{+}(A)$ is divided into the proofs of the following three (independent) propositions:
@@
Let $S\subset\mathbb{R}^{d}$ ($d\geq 3$) with positive reach be a set skewered on $[c,d]$, $c<0<d$, with planar leaves, and $0\in T_{1}(S)$. Then there exists a $\operatorname{B}_{+}$-set $B$ and a multirotation $\rho$ associated with $T_{1}(B)$ such that $(S,0)\sim(\rho(B),0)$.
@@
If $a\in B(b,\tau)$ then (25) is obvious. Assume that $a\in\partial B(b,\tau)$ and denote $k:=\dim\operatorname{Tan}(A,a)$ and $H:=\{x\in\mathbb{R}^{d}:\,\langle x,b-a\rangle>0\}$. By Corollary 2.13 there exists $v\in\operatorname{Tan}(A,a)\cap S^{d-1}$ such that $\angle(v,b-a)<\pi/6$, hence, $v\in H$, thus $\operatorname{Tan}(A,a)\cap H\neq\emptyset$. Then also the relative interior of $\operatorname{Tan}(A,a)$ meets $H$ by [25, Corollary 6.3.2], hence $\operatorname{Tan}(A,a)\cap H$ contains a $k$-dimensional ball. It follows easily from the definition of the tangent cone that $\operatorname{Tan}(A,a)\cap H\subset\operatorname{Tan}(A^{*},a)$, hence $\dim\operatorname{Tan}(A^{*},a)\geq k$. As the other inequality is obvious, we have $\dim\operatorname{Tan}(A^{*},a)=k$ and the proof is finished. ∎
@@
Choose $0<r<{\rm reach}\,A$ and an affine isometry $\Omega:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $\Omega(a)=0$ and $\operatorname{Tan}(\Omega(A),x_{0})=W$. Clearly, $E:=\Omega(A)$ fulfills $r<{\rm reach}\,E$ and $\dim E\leq 2$.
Applying Theorem 4.4 we find that there exists a $1$-dimensional $C^{1,1}$ surface $\Gamma$ associated with $V:=W^{\perp}$ and $\delta>0$ such that
@@
By definition,
$\Gamma=\{w+\varphi(w):\ w\in W\},$
where $\varphi:W\to V$ is a $C^{1,1}$ mapping. Now define the mapping
@@
To this end, first note that obviously $S\subset\pi^{-1}([c,d])$ and we have $0\in T_{1}(S)$ by Lemma 3.2 (iii). Further, consider an arbitrary $w\in[c,d]\subset W$ and set $z:=\Phi^{-1}(w)\in\Gamma$. Then (28) implies $z\in B(0,\delta)$ and so $z\in E$ by (26). Since $w=\Phi(z)$ we have $w\in S$. Thus $c,d]\subset S$. Finally, let $x\in T_{1}(M)$ be given. Using Lemma 5.11 and Lemma 3.2 (iii) we obtain $x\in T_{1}(\Phi(E))=\Phi(T_{1}(E))$. So $x=\Phi(z)$ for some $z\in T_{1}(E)$. Since $z=\Phi^{-1}(x)$ and $x\in S\subset\overline{B}(0,\omega)$, we obtain $z\in B(0,\delta)$ by (28). Thus $z\in\Gamma$ by (26) and consequently $x=\Phi(z)\in W\cap\overline{B}(0,\omega)=[c,d]$. This completes the proof of (29).
@@
$(A,a)\approx(\Omega(A),0)\approx(E,0)\approx(\Phi(E),0)\sim(S,0),$
@@ 5.3. Some properties of skewered sets with positive reach
The following two lemmas will be used in the proof of the most difficult Proposition 5.9.
@@
Let $A\subset\mathbb{R}^{d}$ ($d\geq 2$) be a set with positive reach with $\dim A\leq 2$ skewered on a segment $p,q]\subset\mathbb{R}$, $0<r<{\rm reach}\,A$ and $x_{0}\in T_{1}(A)$. Then the following statements hold.
@@
Assertion (iv) follows from (iii) in the case ${\rm diam}\,A<r$, and from Remark 5.7 (b) in the case of a $B$-set AA. The property (v) follows from (iv) and the closedness of $A$.
@@
where $I$ denotes the identity matrix and $\mathcal{D}=\{s=t_{0}<t_{1}<\dots<t_{k}=t\}$ a partition of $[s,t]$ with mesh $\|\mathcal{D}\|$. We also set $\mu(s,s):=I$, $s\geq 0$. Then $\mu(s,t)$ is multiplicative in the sense that
@@
$\alpha(s,t):=\sum_{s<b_{\lambda}\leq t}\left(\pi_{\psi^{\lambda}(a_{\lambda})}-\pi_{\psi^{\lambda}(b_{\lambda})}\right),\quad 0\leq s\leq t.$
(an empty sum is considered as 0). The sum converges (absolutely) since, due to (38),
@@
Consider arbitrary $x,y\in A$. We shall verify conditions (2) and (3) of Fact 2.3 for some $c>0$.
@@
Case III: $x\in\overline{A}_{\beta}$ and $y\in\overline{A}_{\lambda}$ with $\beta\neq\lambda$, $\beta,\lambda\in\Lambda$. Assume that $b_{\beta}\leq a_{\lambda}$ (the opposite case can be done quite analogously). Since both $b_{\beta}$ and $a_{\lambda}$ belong to $T_{1}(A)$, $|x-b_{\beta}|<r/2$, $|y-a_{\lambda}|<r/2$ and $|x-a_{\lambda}|<r/2$, we have from Lemma 5.12 (iii)
@@
with $c_{3}^{\prime}:=\max\{c_{2}\frac{2}{\sqrt{3}},c_{1}\}$, where we have used (48), hence (2) holds. Also, using the estimates from Case II, we get
@@
Using Lemma 5.12 (ii) we obtain that ${\rm reach}\,A_{1}>0$, $\dim A_{1}\leq 2$ and $A_{1}$ is skewered on $\rho-\delta,\rho+\delta]$. Applying Lemma 2.21 we get also ${\rm reach}\,\tilde{A}>0$, and clearly we have $\dim\tilde{A}\leq 2$. Equations (47) together with $F|_{A}=f$ yields $\rho-\delta,\rho+\delta]\subset\tilde{A}\subset\pi^{-1}([\rho-\delta,\rho+\delta])$, and, with Lemma 3.2 (iii) also
(49) $T_{1}(\tilde{A})=T_{1}(F(A_{1}))=F(T_{1}(A_{1}))=T_{1}(A_{1})\subset[\rho-\delta,\rho+\delta],$
@@
Further, for any $\lambda\in\Lambda$, the leave $A_{\lambda}$ fulfills $F(A_{\lambda})\subset V_{\lambda}$, where $V_{\lambda}\in G(d,2)$ is the (two-dimensional by (41)) image of the linear mapping $\mu_{a_{\lambda}}\circ\pi_{T_{\lambda}}$ (this follows directly from the fact $f=F|_{A}$ and from the definition of $f$). Note that $T_{1}(A_{1})=T_{1}(A)\cap[\rho-\delta,\rho+\delta]$ by Lemma 5.11, hence each leave of $A_{1}$ is a subset of a leave of $A$. Also, each leave of $\tilde{A}$ is the $F$-image of a leave of $A_{1}$ by (49) and (47). Thus we get that $\tilde{A}$ has planar leaves.
@@
In this subsection we will work with multirotations (see Definition 1.2). We will use the easy fact that if $\rho:R^{d}\to\mathbb{R}^{d}$ is a multirotation than it is a bijection and $\rho^{-1}$ is again a multirotation associated with the same set as $\rho$.
@@
We will treat both cases together, using that there exist $E>0$ and $F>0$ such that in both cases, for each $z\in T_{1}(A)$ and $x\in A$,
@@
To prove this claim, first observe that Corollary 2.13 and Proposition 2.12 imply that (50) holds with $E=E_{1}:=2/\sqrt{3}$ and $F=F_{1}:=2/3r$ whenever |zx|<r|z-x|<r and so, in case (a), for all $z\in T_{1}(A)$ and $x\in A$.
In case (b) (in which [c,d]=[r,r][c,d]=[-r,r] or $[c,d]=[0,r]$) recall that functions $\varphi$, $\psi$ from Definition 1.1 ((i) or (ii)) are $L$-Lipchitz for some $L>0$. It is easy to see that $\varphi(z)=\psi(z)=0$ (since $z\in T_{1}(A)$) and $|x-\pi(x)|\leq\max\{\varphi(\pi(x)),-\psi(\pi(x))\}$. Thus we have $|x-\pi(x)|\leq L|\pi(x)-z|$ and consequently $|x-z|\leq\sqrt{1+L^{2}}|\pi(x)-z|$. So, if $|\pi(x)-z|<r/\sqrt{1+L^{2}}$, then |xz|<r|x-z|<r and so (50) holds with $E=E_{2}:=\sqrt{1+L^{2}}$ and $F=F_{1}$. If $|\pi(x)-z|\geq r/\sqrt{1+L^{2}}$, then

### v2

@@ 4. Local structure of $T_{k}(A)$
where (13) and the $L$-Lipschitzness of $\psi_{k}$ on $M$ were used, so (2) and (3) follow with $c:=\max\{L,\frac{1}{2r}\}$. Now, using Fact 2.3, there exists a $C^{1,1}$ mapping $F:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $F(x)=f(x)$ and $DF(x)=\varphi(x)$ whenever $x\in M$. Note that $F(0)=f(0)=0$. Since $DF(0)=\varphi(0)$ is the identity map, we can apply Fact 2.2 to $F$ and find $\omega>\varepsilon>0$ such that the restriction $\Phi|_{B(0,\varepsilon)}$ is a $C^{1,1}$-diffeomorphism. Choose $\varepsilon>\delta>0$ such that $B(0,\delta)\subset\Phi(B(0,\varepsilon))$ and set $G(x):=\pi_{T^{\perp}}(\Phi^{-1}(x))$ for $x\in T\cap B(0,\delta)$. Then $G:T\cap B(0,\delta)\to T^{\perp}$ is a $C^{1,1}$ mapping which has a $C^{1,1}$ extension $\tilde{G}:T\to T^{\perp}$ by Fact 2.4. Set $\Gamma:=\{x+\tilde{G}(x):\ x\in T\}$. Then $\Gamma$ is a $k$-dimensional $C^{1,1}$ g-surface and $T_{k}(A)\cap B(a,\delta)\subset\Gamma$. Indeed, consider an arbitrary $z\in T_{k}(A)\cap B(a,\delta)\subset M$ and denote $x:=\Phi(z)=f(z)=\pi_{T}(z)$. Then $G(x)=\pi_{T^{\perp}}(\Phi^{-1}(x))=\pi_{T^{\perp}}(z)$ and so $z=\pi_{T}(z)+\pi_{T^{\perp}}(z)=x+G(x)\in\Gamma$. ∎
@@
By Theorem 4.1 and Definition 2.5, we can assign to each $a\in T_{k}(A)$ a $\delta_{a}>0$ and a $k$-dimensional $C^{1,1}$ g-surface $\Gamma_{a}$ such that $T_{k}(A)\cap B(a,\delta_{a})\subset\Gamma_{a}$. Since $\mathbb{R}^{d}$ is a separable metric space, there exists a countable set $S\subset T_{k}(A)$ such that $\bigcup_{a\in T_{k}(A)}B(a,\delta_{a})=\bigcup_{a\in S}B(a,\delta_{a})$. Then
@@
This proves (i). Further, we obtain (ii) from (i) using (14) since each $i$-dimensional $C^{1,1}$ g-surface with $1\leq i\leq k$ is clearly contained in a $k$-dimensional $C^{1,1}$ g-surface, and $T_{0}(A)$ is countable by Lemma 3.3 (iv). ∎
@@
By [22, Theorem 7.5], if $\dim A=k$ then $A$ can be locally covered by finitely many “DC surfaces of dimension kk” (which are even semiconcave if k=d1k=d-1, see [22, Theorem 5.9]). These results, however, do not imply Corollary 4.2 (ii) where $C^{1,1}$ surfaces are used.
@@
Without any loss of generality we will suppose that $a=0$ and $e_{1}\in\operatorname{Tan}(A,0)$. Denote $W:=\operatorname{span}{e_{1}}$ and $\pi:=\pi_{W}$.
@@
$\varphi(w):=\gamma\left(\left(G|_{U}\right)^{-1}(w)\right)-w\in W^{\perp},\quad w\in G(U)\subset W,$
$\varphi:G(U)\subset W\to W^{\perp}$ is $C^{1,1}$ smooth and
@@
Choose $0<\delta<\omega$ such that $(-\delta,\delta)\subset G(U)$. Then, by Fact 2.4, $\varphi|_{(-\delta,\delta)}:(-\delta,\delta)\to W^{\perp}$ has an $C^{1,1}$ extension $\tilde{\varphi}:W\to W^{\perp}$, $\Gamma:=\{w+\tilde{\varphi}(w):\,w\in W\}$ is a $1$-dimensional $C^{1,1}$ g-surface and we have
@@ 5.1. Description of the strategy and some auxiliary notions
The difficulty of Theorem 1.3 consists in the proof of its first part (the proof of the second part is easy). It is advantageous to reformulate and prove the first part using an equivalence relation $\approx$ from the following definition.
@@
Let $A\subset\mathbb{R}^{d}$ ($d\geq 3$) have positive reach, $\dim A\leq 2$ and $a\in T_{1}(A)$. Then there exists a $\operatorname{B}$-set $B\subset\mathbb{R}^{2}\subset\mathbb{R}^{d}$ and a multirotation $\rho$ associated with $T_{1}(B)$ such that $(A,a)\approx(\rho(B),0)$.
@@
    There exists a $\operatorname{B}$-set $B_{2}\subset\mathbb{R}^{2}\subset\mathbb{R}^{d}$ and a multirotation $\rho_{2}$ associated with $T_{1}(B_{2})$ such that $(A,a)\approx(\rho_{2}(B_{2}),0)$.
@@
Let $C,C^{\prime}\subset W:=\operatorname{span}\{e_{1}\}=\mathbb{R}\subset\mathbb{R}^{d}$ ($d\geq 3$) be nonempty compact sets such that $C\cap[p,q]=C^{\prime}\cap[p,q]$ for some given $p<q$. Let $\rho$ be a multirotation associated with $C$. Then there exists a multirotation $\rho^{\prime}$ associated with $C^{\prime}$ such that $\rho|_{\pi^{-1}[p,q]}=\rho^{\prime}|_{\pi^{-1}[p,q]}$.
@@
Let $\rho$ be determined by $2$-rotations ($R_{\lambda}$), where $\lambda$ are components of $W\setminus C$. We define $\rho^{\prime}$ as the multirotation determined by $2$-rotations $(R^{\prime}_{\lambda^{\prime}})$, where $\lambda^{\prime}$ are components of $W\setminus C^{\prime}$. We set $R^{\prime}_{\lambda^{\prime}}:={\rm id}$ if $\lambda^{\prime}\cap[p,q]=\emptyset$ and $R^{\prime}_{\lambda^{\prime}}:=R_{\lambda}$ if $\lambda^{\prime}\cap[p,q]\neq\emptyset$, where $\lambda$ is the (clearly unique) component of $W\setminus C$ such that $\lambda\cap[p,q]=\lambda^{\prime}\cap[p,q]$. It is easy to verify that $\rho^{\prime}$ has the desired properties. ∎
@@
Throughout the section we will use the usual identifications $\mathbb{R}^{d}=\mathbb{R}\times\mathbb{R}^{d-1}$ and $\mathbb{R}=\mathbb{R}\times\{0\}=:W\subset\mathbb{R}^{d}$. Further, we denote $\pi:=\pi_{W}:\mathbb{R}^{d}\to W$.
@@
    Any $\operatorname{B}$-set $B\subset\mathbb{R}^{2}\subset\mathbb{R}^{d}$ is skewered (on $[0,r]$ or [r,r][-r,r] for some $r>0$), both in $\mathbb{R}^{2}$ and in $\mathbb{R}^{d}$. Indeed, $B$ has positive reach in $\mathbb{R}^{2}$ by Lemma 2.23, hence also in $\mathbb{R}^{d}$ by Remark 2.15, and it is easy to see that $0\in T_{1}(B)$. Moreover, if $\varphi,\psi$ are the functions from Definition 1.1, we have
@@
The proof of Proposition 5.3 for the case $a\in T_{1}^{+}(A)$ is divided into the proofs of the following three (mutually independent) propositions:
@@
Let $S\subset\mathbb{R}^{d}$ ($d\geq 3$) with positive reach be a set skewered on $[c,d]$, $c<0<d$, with planar leaves, and $0\in T_{1}(S)$. Then there exists a $\operatorname{B}_{+}$-set $B\subset\mathbb{R}^{2}\subset\mathbb{R}^{d}$ and a multirotation $\rho$ associated with $T_{1}(B)$ such that $S\sim_{0}\rho(B)$.
@@
If $a\in B(b,\tau)$ then (25) is obvious. Assume that $a\in\partial B(b,\tau)$ and denote $k:=\dim\operatorname{Tan}(A,a)$ and $H:=\{x\in\mathbb{R}^{d}:\,\langle x,b-a\rangle>0\}$. By Corollary 2.14 there exists $v\in\operatorname{Tan}(A,a)\cap S^{d-1}$ such that $\angle(v,b-a)<\pi/6$, hence $v\in H$ and thus $\operatorname{Tan}(A,a)\cap H\neq\emptyset$. Then also the relative interior of $\operatorname{Tan}(A,a)$ meets $H$ by [24, Corollary 6.3.2], hence $\operatorname{Tan}(A,a)\cap H$ contains a $k$-dimensional ball. It follows easily from the definition of the tangent cone that $\operatorname{Tan}(A,a)\cap H\subset\operatorname{Tan}(A^{*},a)$, hence $\dim\operatorname{Tan}(A^{*},a)\geq k$. As the other inequality is obvious, we have $\dim\operatorname{Tan}(A^{*},a)=k$ and the proof is finished. ∎
@@
Choose $0<r<{\rm reach}\,A$ and an affine isometry $\Omega:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $\Omega(a)=0$ and $\operatorname{Tan}(\Omega(A),0)=W$. Clearly, $E:=\Omega(A)$ fulfills $r<{\rm reach}\,E$ and $\dim E\leq 2$.
Applying Theorem 4.4 we find that there exists a $1$-dimensional $C^{1,1}$ surface $\Gamma$ and $\delta>0$ such that
@@
Moreover, inspecting the proof of Theorem 4.4, we easily see that we can choose
$\Gamma=\{w+\varphi(w):\ w\in W\}$
with a suitable $C^{1,1}$ mapping $\varphi:W\to W^{\perp}$. Now define the mapping
@@
To this end, first note that obviously $S\subset\pi^{-1}([c,d])$ and we have $0\in T_{1}(S)$ by Lemma 3.2 (iii). Further, consider an arbitrary $w\in[c,d]\subset W$ and set $z:=\Phi^{-1}(w)\in\Gamma$. Then (28) implies $z\in B(0,\delta)$ and so $z\in E$ by (26). Since $w=\Phi(z)$ we have $w\in S$. Thus $c,d]\subset S$. Finally, let $x\in T_{1}(S)$ be given. Using Lemma 5.11 and Lemma 3.2 (iii) we obtain $x\in T_{1}(\Phi(E))=\Phi(T_{1}(E))$. So $x=\Phi(z)$ for some $z\in T_{1}(E)$. Since $z=\Phi^{-1}(x)$ and $x\in S\subset\overline{B}(0,\omega)$, we obtain $z\in B(0,\delta)$ by (28). Thus $z\in\Gamma$ by (26) and consequently $x=\Phi(z)\in W\cap\overline{B}(0,\omega)=[c,d]$. This completes the proof of (29).
@@
$(A,a)\approx(\Omega(A),0)\approx(E,0)\approx(\Phi(E),0)\approx(S,0),$
@@ 5.3. Some properties of skewered sets with positive reach
The following two lemmas will be used in the proof of the most difficult Proposition 5.9. Recall that Notation 5.5 is used.
@@
Let $A\subset\mathbb{R}^{d}$ ($d\geq 2$) be a set with positive reach with $\dim A\leq 2$ skewered on a segment $p,q]\subset\mathbb{R}=W$, $0<r<{\rm reach}\,A$ and $x_{0}\in T_{1}(A)$. Then the following statements hold.
@@
Assertion (iv) follows from (iii) in the case ${\rm diam}\,A<r$, and from Remark 5.7 (b) in the case of a $B$-set AA. Property (v) follows from (iv) and the closedness of $A$.
@@
where $I$ denotes the identity matrix and $\mathcal{D}=\{s=t_{0}<t_{1}<\dots<t_{k}=t\}$ is a partition of $[s,t]$ with mesh $\|\mathcal{D}\|$. We also set $\mu(s,s):=I$, $s\geq 0$. Then $\mu(s,t)$ is multiplicative in the sense that
@@
$\alpha(s,t):=\sum_{s<b_{\lambda}\leq t}\left(\pi_{\psi^{\lambda}(a_{\lambda})}-\pi_{\psi^{\lambda}(b_{\lambda})}\right),\quad 0\leq s\leq t$
(an empty sum is defined as 0). The sum converges (absolutely) since, due to (38),
@@
Now we are going to extend $f$ using Fact 2.3. To this end, we will find $c>0$ such that conditions (2) and (3) hold for $f$, $\varphi$ and all $x,y\in A$. We will distinguish several cases.
@@
Case III: $x\in\overline{A}_{\beta}$ and $y\in\overline{A}_{\lambda}$ with $\beta\neq\lambda$, $\beta,\lambda\in\Lambda$. Assume that $b_{\beta}\leq a_{\lambda}$ (the case $b_{\lambda}\leq a_{\beta}$ can be done quite analogously). Since both $b_{\beta}$ and $a_{\lambda}$ belong to $T_{1}(A)$, $|x-b_{\beta}|<r$, $|y-a_{\lambda}|<r$ and $|x-a_{\lambda}|<r$, we have from Lemma 5.12 (iii)
@@
with $c_{3}^{\prime}:=\max\{c_{2}\frac{2}{\sqrt{3}},c_{1}\}$, where we have used (48). Also, using the estimates from Case II, we get
@@
Using Lemma 5.12 (ii) we obtain that ${\rm reach}\,A_{1}>0$, $\dim A_{1}\leq 2$ and $A_{1}$ is skewered on $\rho-\delta,\rho+\delta]$. Applying Lemma 2.22 we get also ${\rm reach}\,\tilde{A}>0$, and clearly we have $\dim\tilde{A}\leq 2$. Equations (47) together with $F|_{A}=f$ yield $\rho-\delta,\rho+\delta]\subset\tilde{A}\subset\pi^{-1}([\rho-\delta,\rho+\delta])$, and, with Lemma 3.2 (iii) also
(49) $T_{1}(\tilde{A})=T_{1}(F(A_{1}))=F(T_{1}(A_{1}))=T_{1}(A_{1})\subset[\rho-\delta,\rho+\delta];$
@@
Further, for any $\lambda\in\Lambda$, the leave $A_{\lambda}$ fulfills $F(A_{\lambda})\subset V_{\lambda}$, where $V_{\lambda}$ is the image of the linear mapping $\mu_{a_{\lambda}}\circ\pi_{T_{\lambda}}$ (this follows directly from the fact $f=F|_{A}$ and from the definition of $f$). We have $V_{\lambda}\in G(d,2)$ by (41). Note that $T_{1}(A_{1})=T_{1}(A)\cap[\rho-\delta,\rho+\delta]$ by Lemma 5.11, hence each leave of $A_{1}$ is a subset of a leave of $A$. Also, each leave of $\tilde{A}$ is the $F$-image of a leave of $A_{1}$ by (49) and (47). Thus we get that $\tilde{A}$ has planar leaves.
@@
In this subsection we will work with multirotations (see Definition 1.2). We will use the easy fact that if $\rho:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a multirotation than it is a bijection and $\rho^{-1}$ is again a multirotation associated with the same set as $\rho$.
@@
We will treat both cases (a) and (b) together, using that there exist $E>0$ and $F>0$ such that in both cases, for each $z\in T_{1}(A)$ and $x\in A$,
@@
To prove this claim, fix arbitrary $z\in T_{1}(A)$ and $x\in A$. First observe that Corollary 2.14 and Proposition 2.13 imply that (50) holds with $E=E_{1}:=2/\sqrt{3}$ and $F=F_{1}:=2/3r$ whenever |zx|<r|z-x|<r, and so always in case (a).
In case (b) (in which [c,d]=[r,r][c,d]=[-r,r] or $[c,d]=[0,r]$) recall that functions $\varphi$, $\psi$ from Definition 1.1 ((i) or (ii)) are $L$-Lipchitz for some $L>0$. Since $z\in T_{1}(A)$ we have $\varphi(z)=\psi(z)=0$ by Remark 5.7 (b) and clearly $|x-\pi(x)|\leq\max\{\varphi(\pi(x)),-\psi(\pi(x))\}$. Thus we have $|x-\pi(x)|\leq L|\pi(x)-z|$ and consequently $|x-z|\leq\sqrt{1+L^{2}}|\pi(x)-z|$. So, if $|\pi(x)-z|<r/\sqrt{1+L^{2}}$, then |xz|<r|x-z|<r and so (50) holds with $E=E_{2}:=\sqrt{1+L^{2}}$ and $F=F_{1}$. If $|\pi(x)-z|\geq r/\sqrt{1+L^{2}}$, then